OpenStudy (anonymous):

Solve for x. 2x^2 - 4x = 0 Select one of the options below as your answer: A. 0, -4 B. 0, -2 C. 0, 2 D. 2, 4

OpenStudy (anonymous):

0,2

OpenStudy (anonymous):

simplify to x(2x-4)=0. Then this equals zero when x = 0 and when 2x - 4 = 0. (x=2).

OpenStudy (anonymous):

ohkay thanks